# To Err Is Human, but Llamas Can Learn It Too Анотацiя

### v1

@@ 4.2 Artificial Error Generation with Smaller Models
Table 3 shows the results of the analysis. Both models can learn valuable information that improves performance beyond what is achieved with fine-tuning on gold data alone. Notably, errors generated by the NLLB model are particularly effective, delivering results close to those achieved by LM-generated errors in Estonian and German, almost matching the performance of LLama-based models. However, for Ukrainian, NLLB-generated errors fall behind probabilistic reverse-speller errors. The Ukrainian NLLB zero-shot GEC performance is also significantly lower than for Estonian or German (see more in Appendix C) or English that Luhtaru et al. (2024) also tested.
@@
We can also see that the scores before gold fine-tuning highlight that Ukrainian scores are notably low across all methods. However, these scores recover well after fine-tuning, suggesting the synthetic data may not align well with the text domain or error types specific to the Ukrainian language. Estonian and German models show higher scores for models trained with just AEG data and improve less drastically with fine-tuning.
@@ 4.3 Artificial Error Generation with Prompting
To assess the capability of generating errors without additional LM training, we utilize advanced commercial models, specifically exploring the efficiency of error generation through prompting GPT-3.5-turbo with datasets comprising 100,000 sentences. We later also explore the effectiveness of GPT-4-turbo in a more limited setting (see Section 4.4).
The generation cost depends on the sum of input and completion tokens. Ukrainian, our most expensive language, had the highest number of tokens per 100,000 sentences: 98 million input and 12 million completion tokens. The cost for input tokens with GPT-3.5-Turbo in USD is $147, and for completion tokens, it is $25 – in total, $172 for generating 100,000 Ukrainian sentences. In comparison, the costs with GPT-4-Turbo would have been $983 and $370, respectively******openai.com/pricing.
Table 4 shows the results of continued pre-training Llama-based models on the same amount of sentences (100,000) with synthetic errors from prompting or fine-tuning. In terms of error correction quality after gold fine-tuning, employing GPT-3.5-turbo for prompting and fine-tuning Llama-2-based models are both viable strategies for artificial error generation, as they lead to very close F scores in all three languages (with a slight difference in favor of fine-tuning errors for German: 75.58 vs 76.25).
@@
When comparing LLama model scores for 100k to the ones with only gold tuning (see Table 2), we can see that although scores increase more modestly, only 100k examples of synthetic data increase the scores more for German (almost 2 F-score points), a bit for Estonian (around 0.4 points) and stay the same for Ukrainian with higher precision and lower recall. This shows a possible text domain mismatch between the human error train/test data and our choice of monolingual sentences. This negative effect is alleviated with higher numbers of pretraining AEG data in the 1M sentence experiments.
4.4 Error Generation Quality
@@
Our findings indicate that the precision of all synthetic data closely matches that of high-quality (gold) data in both Estonian and German, as illustrated in Figure 1. A notable distinction, however, is observed in recall rates. For Estonian and German, the recall for errors generated by LMs is more comparable to human-generated (gold) data than errors produced through probabilistic methods.
@@
Ukrainian scores with synthetic data are substantially worse than gold data, regardless of the AEG method. Still, recall for LM-generated errors is significantly higher than for simple probabilistic errors. This might be due to a larger mismatch in the text domain or error frequency. Ukrainian UA-GEC data predominantly contains punctuation errors (43%) and has a two times smaller error rate than German (8.2 vs 16.8) Syvokon et al. (2023).
@@
Comparing GPT-3.5-turbo with GPT-4-turbo, we find similar performance overall. However, for Estonian, GPT-4-turbo exhibits higher recall but lower precision. For German, GPT-4-turbo shows reductions in both precision and recall. Performance is nearly identical for Ukrainian between the two models. Overall, the F scores of GPT-4-turbo are slightly lower for Estonian and German and marginally higher for Ukrainian compared to GPT-3.5.
@@
When analyzing the recall for various error categories in Estonian, it is evident that our models trained with AEG data particularly face challenges in inserting missing punctuation marks and correcting errors related to word order, as depicted in Figure 2. Errors generated probabilistically excel in identifying spelling mistakes and can correct certain errors in noun and verb forms. However, they generally perform poorly in addressing issues beyond spelling errors. This comparison suggests that our learned and prompted synthetic errors are much more similar to naturally made human errors.
5 Discussion
We investigated contemporary methods for generating artificial errors (AEG) for Estonian, German, and Ukrainian languages with relatively scarce resources. These languages have approximately 10k, 20k, and 30k error correction examples derived from corpora with varied error type distributions. The Estonian and Ukrainian corpora notably include language learner texts, characterized by a high frequency of errors, whereas the Ukrainian corpus contains more native-speaker texts with less edits.
Across these languages, our primary approach (AEG using Llama-based models and grammar error correction with Llama-based language models) demonstrated consistent efficacy after fine-tuning with error correction examples. This success underscores the value of the learned error generation method over the probabilistic reverse-speller approach, as evidenced by improved precision and recall based on reference metrics. The other methods – prompting and smaller models – also consistently prove useful.
However, before fine-tuning with gold-standard GEC examples, we observed divergent language behaviors, raising questions about potential overfitting to these test sets and the generalizability of methods trained on specific datasets. For instance, our Ukrainian test set presented challenges for all methods lacking specific training data, including those involving GPT-4 models. It remains unclear whether methods that tend towards paraphrasing and fluency edits, including GPT models Coyne et al. (2023), fail to align with the precise edits needed, overcorrect, or generate incorrect corrections. Critiques of current GEC metrics, which are argued to poorly correlate with human judgments Sakaguchi et al. (2016); Östling et al. (2023), suggest that true quality assessment may require human evaluation — a step beyond the scope of our study.
The observed performance with generated errors may also relate to mismatches between the chosen monolingual sentences and the original GEC human data. While our study utilizes web texts, resembling the essay-like texts typically employed in GEC and differing from native speaker constructions, these sentences might be simpler than those the model is accustomed to handling.
@@ 6 Conclusion
In conclusion, our research demonstrates the significant potential of Llama-based LMs in addressing the challenges of GEC for low-resource languages. We have successfully developed state-of-the-art systems for Estonian, Ukrainian, and German by leveraging these models as both correctors and synthetic data generators. We also explore other methods for AEG and show that prompting stronger commercial LMs is another way of generating high-quality data, and fine-tuning smaller models also has potential when the resources are more limited.
Potential directions for future work include tuning LMs to perform AEG and GEC multilingually (like Rothe et al., 2021; Luhtaru et al., 2024), applying our proposed AEG methods to monolingual data of a more similar text domain to the benchmarks Oda (2023). An interesting direction would be to test these methods with high-resource GEC languages (English, Chinese).
@@ 7 Limitations
Furthermore, our study lacks human evaluation, a component for more reliably assessing the real-world efficacy of GEC systems.
@@
- Coyne et al. (2023) Steven Coyne, Keisuke Sakaguchi, Diana Galvan-Sosa, Michael Zock, and Kentaro Inui. 2023. Analyzing the performance of gpt-3.5 and gpt-4 in grammatical error correction.
@@
- Fang et al. (2023) Tao Fang, Xuebo Liu, Derek F. Wong, Runzhe Zhan, Liang Ding, Lidia S. Chao, Dacheng Tao, and Min Zhang. 2023. TransGEC: Improving grammatical error correction with translationese. In Findings of the Association for Computational Linguistics: ACL 2023, pages 3614–3633, Toronto, Canada.
@@
- Oda (2023) Mikio Oda. 2023. Training for grammatical error correction without human-annotated L2 learners’ corpora. In Proceedings of the 18th Workshop on Innovative Use of NLP for Building Educational Applications (BEA 2023), pages 455–465, Toronto, Canada.
@@
- Sakaguchi et al. (2016) Keisuke Sakaguchi, Courtney Napoles, Matt Post, and Joel Tetreault. 2016. Reassessing the Goals of Grammatical Error Correction: Fluency Instead of Grammaticality. Transactions of the Association for Computational Linguistics, 4:169–182.
@@
- Tang et al. (2021) Yuqing Tang, Chau Tran, Xian Li, Peng-Jen Chen, Naman Goyal, Vishrav Chaudhary, Jiatao Gu, and Angela Fan. 2021. Multilingual translation from denoising pre-training. In Findings of the Association for Computational Linguistics: ACL-IJCNLP 2021, pages 3450–3466, Online.
@@
- Touvron et al. (2023) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, Dan Bikel, Lukas Blecher, Cristian Canton Ferrer, Moya Chen, Guillem Cucurull, David Esiobu, Jude Fernandes, Jeremy Fu, Wenyin Fu, Brian Fuller, Cynthia Gao, Vedanuj Goswami, Naman Goyal, Anthony Hartshorn, Saghar Hosseini, Rui Hou, Hakan Inan, Marcin Kardas, Viktor Kerkez, Madian Khabsa, Isabel Kloumann, Artem Korenev, Punit Singh Koura, Marie-Anne Lachaux, Thibaut Lavril, Jenya Lee, Diana Liskovich, Yinghai Lu, Yuning Mao, Xavier Martinet, Todor Mihaylov, Pushkar Mishra, Igor Molybog, Yixin Nie, Andrew Poulton, Jeremy Reizenstein, Rashi Rungta, Kalyan Saladi, Alan Schelten, Ruan Silva, Eric Michael Smith, Ranjan Subramanian, Xiaoqing Ellen Tan, Binh Tang, Ross Taylor, Adina Williams, Jian Xiang Kuan, Puxin Xu, Zheng Yan, Iliyan Zarov, Yuchen Zhang, Angela Fan, Melanie Kambadur, Sharan Narang, Aurelien Rodriguez, Robert Stojnic, Sergey Edunov, and Thomas Scialom. 2023. Llama 2: Open foundation and fine-tuned chat models. arXiv preprint arXiv:2307.09288.
@@
- Östling et al. (2023) Robert Östling, Katarina Gillholm, Murathan Kurfalı, Marie Mattson, and Mats Wirén. 2023. Evaluation of really good grammatical error correction.
Додаток A Prompts
We present the prompts used to generate 1) 100,000 sets with GPT-3.5-Turbo and 2) preliminary sets with GPT-4-Turbo in Tables 5, 6, 7 for Estonian, German, and Ukrainian respectively.
@@
Додаток B Training details
B.1 Llama-based models
The models are trained on 4 AMD MI250x GPUs (each acting as 2 GPUs).
| Parameter | Value |
| LR | 5e-6 |
| LRfinal | 5e-7 |
| LR-schedule | linear |
| Epochs | 3 |
| Max sequence length | 1024 |
| Batch size (total) | 128 |
| Gradient clipping | 1.0 |
| Weight decay | 0.1 |
| Optimizer | AdamW |
| Precision | bf16 |
| DeepSpeed | Zero Stage 2 |
| Parameter | Value |
| LR | 2e-5 |
| LRfinal | 2e-6 |
| LR-schedule | linear |
| Updates | 19080 |
| Max sequence length | 1024 |
| Batch size (total) | 256 |
| Gradient clipping | 1.0 |
| Weight decay | 0.1 |
| Optimizer | AdamW |
| Precision | bf16 |
| DeepSpeed | Zero Stage 2 |
B.2 NLLB-based models
We follow the training process specified by Luhtaru et al. (2024), including hyperparameters. The training is conducted on an AMD MI250x GPU. We are training the AEG models for 20 epochs and picking the 15th after arbitrary manual evaluation and testing sets on checkpoints 5, 10, 15, and 20. The data for NLLB models is first normalized with Moses script††††††https://github.com/pluiez/NLLB-inference/blob/main/preprocess/normalize-punctuation.perl, and we use the SentencePiece model Kudo and Richardson (2018) for untokenized text.
B.3 mT5-based models
Додаток C NLLB correction
| Lang | Zero-shot | Gold fine-tuning | ||||
| P | R | F | P | R | F | |
| Estonian | 43.89 | 45.31 | 44.17 | 61.14 | 49.48 | 58.39 |
| Ukrainian | 8.24 | 31.57 | 9.67 | 35.62 | 34.1 | 35.31 |
| German | 43.66 | 41.52 | 43.22 | 73.71 | 67.75 | 72.44 |

### v2

@@ 4.2 Artificial Error Generation with Smaller Models
Table 3 shows the results of the analysis. Both models can learn valuable information that improves performance beyond what is achieved with fine-tuning on gold data alone. Notably, errors generated by the NLLB model are particularly effective, delivering results close to those achieved by LLM-generated errors in Estonian and German, almost matching the performance of LLama-based models. However, for Ukrainian, NLLB-generated errors fall behind probabilistic reverse-speller errors. This is likely because the dataset contains many special punctuation characters that get normalized during preprocessing (see more in Appendix C).
@@
We can also see that the scores before gold fine-tuning highlight that Ukrainian scores are notably low across all methods. However, these scores recover well after fine-tuning, suggesting the synthetic data may not align well with the text domain or error types specific to the Ukrainian language. This can be related to the unique characteristics of the Ukrainian dataset, which also causes GPT-4 to struggle with the GEC task. Estonian and German models show higher scores for models trained with just AEG data and improve less drastically with fine-tuning.
@@ 4.3 Artificial Error Generation with Prompting
To assess the capability of generating errors without additional LLM training, we utilize advanced commercial models, specifically exploring the efficiency of error generation through prompting GPT-3.5-turbo with datasets comprising 100,000 sentences. We later also explore the effectiveness of GPT-4-Turbo in a more limited setting (see Section 4.4).
The generation cost depends on the sum of input and completion tokens. Ukrainian, our most expensive language, had the highest number of tokens per 100,000 sentences: 98 million input and 12 million completion tokens. The cost for input tokens with GPT-3.5-Turbo in USD is $147, and for completion tokens, it is $25 – in total, $172 for generating 100,000 Ukrainian sentences. In comparison, the costs with GPT-4-Turbo would have been $983 and $370, respectively‡‡‡‡‡‡https://openai.com/pricing.
Table 4 shows the results of continued pre-training Llama-based models on the same amount of sentences (100,000) with synthetic errors from prompting or fine-tuning. In terms of error correction quality after gold fine-tuning, employing GPT-3.5-turbo for prompting and fine-tuning Llama-2-based models are both viable strategies for AEG, as they lead to very close F0.5 scores in all three languages (with a slight difference in favor of fine-tuning errors for German: 75.58 vs 76.25).
@@
When comparing LLama model scores for 100k to the ones with only gold tuning (see Table 2), we can see that although scores increase more modestly, only 100k examples of synthetic data increase the scores more for German (almost 2 F0.5-score points), a bit for Estonian (around 0.4 points) and stay the same for Ukrainian with higher precision and lower recall. The scores for models trained with 100k sentences are mostly lower than those trained with 1M reverse-speller errors, which indicates that the data quantity jump from 100,000 to 1M plays a significant role.
4.4 Quality Compared to Human Data
@@
Our findings indicate that the precision of all synthetic data closely matches that of high-quality (gold) data in both Estonian and German, as illustrated in Figure 1. A notable distinction, however, is observed in recall rates. For Estonian and German, the recall for errors generated by LLMs is more comparable to human-generated (gold) data than errors produced through probabilistic methods.
@@
Ukrainian scores with synthetic data are substantially worse than gold data, regardless of the AEG method. Still, recall for LLM-generated errors is significantly higher than for simple probabilistic errors. This might be due to a larger mismatch in the text domain or error frequency. The dataset is heavily composed of punctuation errors, which could be more challenging for LLMs to generate, as they have been shown to struggle with correcting such errors Katinskaia and Yangarber (2024).
@@
Comparing GPT-3.5-Turbo with GPT-4-Turbo, we find similar performance overall. However, for Estonian, GPT-4-Turbo exhibits higher recall but lower precision. For German, GPT-4-Turbo shows reductions in both precision and recall. Performance is nearly identical for Ukrainian between the two models. Overall, the F0.5 scores of GPT-4-Turbo are slightly lower for Estonian and German and around the same for Ukrainian compared to GPT-3.5-Turbo.
@@
When analyzing the recall for various error categories in Estonian, it is evident that our models trained with AEG data particularly face challenges in inserting missing punctuation marks and correcting errors related to word order, as depicted in Figure 2. Errors generated probabilistically excel in identifying spelling mistakes and can correct certain errors in noun and verb forms. However, they generally perform poorly in addressing issues beyond spelling errors.
4.5 Evaluation of Generated Errors: Case Study with Estonian
We labeled 100 LLM-generated sentences from different sets to determine if the errors made by models are similar to those in the training corpus.
Based on the annotations, we can categorize a large proportion of the changes according to the annotation scheme, but there is still a considerable amount of problematic edits (2̃5-45%) (see Figure 3 and Table 7 in Appendix B). The human evaluation also indicates that the models differ in their error rates. GPT models generate fewer problematic errors overall, but the error category distribution seems more similar to human data with Llama-based models. This is likely due to a fine-tuning approach instead of prompting.
As mentioned in the last section, compared to human data, all models trained with generated data, correct far fewer word order and missing punctuation errors, and lexical changes are not well corrected either. These results can be partially explained by examining the different error types in generated data, where the same types are not as well represented as in human data. Most problematic edits involve generating lexical errors, which often were synonymous or changed the original meaning of the sentence, which could explain the poor performance in correcting lexical errors. On the other hand, verb or nominal form and spelling errors were better or almost as well corrected as by a model trained with gold data, and the data contained more errors in these categories. This shows that correction recall is closely tied to the error types present in the training data, and the data generated with our approach generates realistic error types that help correction in these categories.
@@ 5 Conclusion
In conclusion, our research demonstrates the significant potential of Llama-based LLMs in addressing the challenges of GEC for low-resource settings. We have successfully developed state-of-the-art systems for Estonian, Ukrainian, and German by leveraging these models as both correctors and synthetic data generators. We also explore other methods for AEG and show that prompting stronger commercial LLMs is another way of generating high-quality data, and fine-tuning smaller models also has potential when the resources are more limited.
@@ 6 Limitations
Furthermore, our study lacks human evaluation of GEC systems, a component for more reliably assessing the real-world efficacy of GEC systems.
7 Acknowledgements
This work was partially supported by the Estonian Research Council grant PRG2006 (Language Technology for Low-Resource Finno-Ugric Languages and Dialects) as well as the Institute of the Estonian Language grant LLTAT24472 (Autocorrect for Estonian as a 2nd language for students and teachers).
We acknowledge the EuroHPC Joint Undertaking for awarding this project access to the EuroHPC supercomputer LUMI, hosted by CSC (Finland) and the LUMI consortium through a EuroHPC Regular Access call.
@@ 7 Acknowledgements
- Allkivi-Metsoja et al. (2022) Kais Allkivi-Metsoja, Karina Kert Kaisa Norak, Silvia Maine, and Pille Eslon. 2022. Error classification and annotation of learner language for developing estonian grammar correction.
@@
- Coyne et al. (2023a) Steven Coyne, Keisuke Sakaguchi, Diana Galvan-Sosa, Michael Zock, and Kentaro Inui. 2023a. Analyzing the performance of gpt-3.5 and gpt-4 in grammatical error correction.
- Coyne et al. (2023b) Steven Coyne, Keisuke Sakaguchi, Diana Galvan-Sosa, Michael Zock, and Kentaro Inui. 2023b. Analyzing the performance of gpt-3.5 and gpt-4 in grammatical error correction.
@@
- Fang et al. (2023a) Tao Fang, Jinpeng Hu, Derek F. Wong, Xiang Wan, Lidia S. Chao, and Tsung-Hui Chang. 2023a. Improving grammatical error correction with multimodal feature integration. In Findings of the Association for Computational Linguistics: ACL 2023, pages 9328–9344, Toronto, Canada.
- Fang et al. (2023b) Tao Fang, Shu Yang, Kaixin Lan, Derek F. Wong, Jinpeng Hu, Lidia S. Chao, and Yue Zhang. 2023b. Is chatgpt a highly fluent grammatical error correction system? a comprehensive evaluation.
@@
- Katinskaia and Yangarber (2024) Anisia Katinskaia and Roman Yangarber. 2024. GPT-3.5 for grammatical error correction. In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), pages 7831–7843, Torino, Italia. ELRA and ICCL.
@@
- Koehn (2004) Philipp Koehn. 2004. Statistical significance tests for machine translation evaluation. In Proceedings of the 2004 Conference on Empirical Methods in Natural Language Processing, pages 388–395, Barcelona, Spain.
@@
- Kuulmets et al. (2024) Hele-Andra Kuulmets, Taido Purason, Agnes Luhtaru, and Mark Fishel. 2024. Teaching llama a new language through cross-lingual knowledge transfer. In Findings of the Association for Computational Linguistics: NAACL 2024, pages 3309–3325, Mexico City, Mexico. Association for Computational Linguistics.
@@
- Touvron et al. (2023a) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, Aurelien Rodriguez, Armand Joulin, Edouard Grave, and Guillaume Lample. 2023a. Llama: Open and efficient foundation language models.
- Touvron et al. (2023b) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, Dan Bikel, Lukas Blecher, Cristian Canton Ferrer, Moya Chen, Guillem Cucurull, David Esiobu, Jude Fernandes, Jeremy Fu, Wenyin Fu, Brian Fuller, Cynthia Gao, Vedanuj Goswami, Naman Goyal, Anthony Hartshorn, Saghar Hosseini, Rui Hou, Hakan Inan, Marcin Kardas, Viktor Kerkez, Madian Khabsa, Isabel Kloumann, Artem Korenev, Punit Singh Koura, Marie-Anne Lachaux, Thibaut Lavril, Jenya Lee, Diana Liskovich, Yinghai Lu, Yuning Mao, Xavier Martinet, Todor Mihaylov, Pushkar Mishra, Igor Molybog, Yixin Nie, Andrew Poulton, Jeremy Reizenstein, Rashi Rungta, Kalyan Saladi, Alan Schelten, Ruan Silva, Eric Michael Smith, Ranjan Subramanian, Xiaoqing Ellen Tan, Binh Tang, Ross Taylor, Adina Williams, Jian Xiang Kuan, Puxin Xu, Zheng Yan, Iliyan Zarov, Yuchen Zhang, Angela Fan, Melanie Kambadur, Sharan Narang, Aurelien Rodriguez, Robert Stojnic, Sergey Edunov, and Thomas Scialom. 2023b. Llama 2: Open foundation and fine-tuned chat models. arXiv preprint arXiv:2307.09288.
@@
- Wu et al. (2023) Haoran Wu, Wenxuan Wang, Yuxuan Wan, Wenxiang Jiao, and Michael Lyu. 2023. Chatgpt or grammarly? evaluating chatgpt on grammatical error correction benchmark.
@@
- Zhang et al. (2023) Yue Zhang, Leyang Cui, Deng Cai, Xinting Huang, Tao Fang, and Wei Bi. 2023. Multi-task instruction tuning of llama for specific scenarios: A preliminary study on writing assistance.
@@
Додаток A Training details
A.1 Llama-based models
The models are trained on 4 AMD MI250x GPUs (each acting as 2 GPUs).
| Parameter | Value |
| LR | 5e-6 |
| LRfinal | 5e-7 |
| LR-schedule | linear |
| Epochs | 3 |
| Max sequence length | 1024 |
| Batch size (total) | 128 |
| Gradient clipping | 1.0 |
| Weight decay | 0.1 |
| Optimizer | AdamW |
| Precision | bf16 |
| DeepSpeed | Zero Stage 2 |
| Parameter | Value |
| LR | 2e-5 |
| LRfinal | 2e-6 |
| LR-schedule | linear |
| Updates | 19080 |
| Max sequence length | 1024 |
| Batch size (total) | 256 |
| Gradient clipping | 1.0 |
| Weight decay | 0.1 |
| Optimizer | AdamW |
| Precision | bf16 |
| DeepSpeed | Zero Stage 2 |
A.2 NLLB-based models
We follow the training process specified by Luhtaru et al. (2024), including hyperparameters. The training is conducted on an AMD MI250x GPU. We are training the AEG models for 20 epochs and picking the 15th after arbitrary manual evaluation and testing sets on checkpoints 5, 10, 15, and 20. The data for NLLB models is first normalized with Moses script, and we use the SentencePiece model Kudo and Richardson (2018) for untokenized text.
A.3 mT5-based models
Додаток B Problematic edits
We further explore the human annotation results discussed in section 4.5. Table 7 displays the percentage of problematic error types out of all errors generated by the model.
| Type | Llama | GPT-3.5 | GPT-4 |
| O | 10.83 | 4.71 | 9.07 |
| HALL | 22.72 | 11.11 | 3.75 |
| SYN | 6.16 | 6.4 | 7.5 |
| INACC | 2.12 | 5.39 | 1.38 |
| UNREC | 3.82 | 6.73 | 3.94 |
| Total % | 45.65 | 34.34 | 25.64 |
Додаток C NLLB correction
| Lang | P | R | F0.5 |
| Estonian | 43.89 | 45.31 | 44.17 |
| Ukrainian | 8.24 | 31.57 | 9.67 |
| German | 43.66 | 41.52 | 43.22 |
Додаток D Prompts
We present the prompts used to generate 1) 100,000 sets with GPT-3.5-Turbo and 2) preliminary sets with GPT-4-Turbo in Tables 9, 10, 11 for Estonian, German, and Ukrainian respectively.